find and explain what error led to the conclusion that 1=2?
true a=b
true (a)(b)=(b)(b)
true ab=b^2
true ab-a^2=b^2-a^2
true a(b-a)=(b-a)(b+a)
false a(b-a)/(b-a)=(b-a)(b+a)/(b-a) is it undefined?
false a=b+a
false a=a+a
false a=2a
false 1=2

Answer by Edwin McCravy(20060)

find and explain what error led to the conclusion that 1=2?
true a=b
true (a)(b)=(b)(b)
true ab=b^2
true ab-a^2=b^2-a^2
true a(b-a)=(b-a)(b+a)
false a(b-a)/(b-a)=(b-a)(b+a)/(b-a) is it undefined?

Yes that's undefined because the first step is a=b, so a-b=0,
and division by 0 is impossible, so the rest is not valid.

false a=b+a
false a=a+a
false a=2a
false 1=2

Edwin